a card is drawn at random from a 52 card deck. tell whether event A and B are inclusive or mutually exclusive. Then find p(A or B). please
1.A; The card is a number less than 5.
B; The card is a jack, king or queen.
2.A; the card is black
B; the card is a number greater then 4

3.A; The card is red
B; The card is the ace of spades.
4.A;The card is a 2 or 3
B;The card is not a heart.
5.A; The card is a diamond
B; the card is a spade.???

Answer by stanbon(57967)

a card is drawn at random from a 52 card deck.
tell whether event A and B are inclusive or mutually exclusive.
Then find p(A or B). please
1.A; The card is a number less than 5.
B; The card is a jack, king or queen.
---
Mutually Exclusive
P(A or B) = P(A)+P(B)= P(2,3,4)+P(j,k,q) = [12/52]+[12/52] = 24/52
-----------------------

2.A; the card is black
B; the card is a number greater then 4
Inclusive:
P(A or B) = P(A)+P(B)-P(AUB) = [26+40-20]/52 = 46/52
-----------------------
Cheers,
Stan H.
==========
Can you handle the rest?
----
3.A; The card is red
B; The card is the ace of spades.
4.A;The card is a 2 or 3
B;The card is not a heart.
5.A; The card is a diamond
B; the card is a spade.???